OpenStudy (lgbasallote):

Eight horses raced to reach the finish line. You picked 3 horses at random to land in either first or second place (or even both) for your racing ticket. If none of the three ended in first or second, you lose. What is the probability that your ticket wins if the race result is random?

OpenStudy (anonymous):

I think this one was already done

OpenStudy (lgbasallote):

why is the solution to this \[\frac{3C1 \times 5C2 + 3C2}{8C2}\] why not \[\frac{2(3C1 \times 5C2) + 3C2}{8C2}\]

OpenStudy (lgbasallote):

@sauravshakya last time i asked for a solution...now im asking why my solution is wrong

OpenStudy (anonymous):

Yeah

OpenStudy (lgbasallote):

3C1 * 5C2 should be interchangeable right?

OpenStudy (anonymous):

Sorry but I did